OpenStudy (anonymous):

Find the 20th term of the arithmetic sequence 7, 3, -1, -5, … I got -62 but thats not even a choice

rishavraj (rishavraj):

hey the first term is 7 and the common differnce is -4...so just apply the formula to get the 20th term...i guess u might have done calculation mistake

OpenStudy (anonymous):

what is the formula again?

OpenStudy (anonymous):

oh i got it its -69

rishavraj (rishavraj):

its --> a + (n - 1)*d a = first term d= common difference n= no of term u wanna find the value of here n= 20...

rishavraj (rishavraj):

@Ratedover u r correct/...

OpenStudy (anonymous):

Find S15 for 1 + 1.5 + 2.25 + 3.375 +... I know the formula

rishavraj (rishavraj):

S15 does it mean sum of 15 terms???

OpenStudy (anonymous):

yes and i got it lol thanks one more question Find the 10th term of the sequence whose nth term is given by the formula tn = (-1)n+1/n. I have no clue

rishavraj (rishavraj):

isnt it given wether its ap or gp???

OpenStudy (anonymous):

what..

rishavraj (rishavraj):

its damn too easy just plug 10 on place of "n"

rishavraj (rishavraj):

@Ratedover solved??

OpenStudy (anonymous):

oh.... well i feel stupid

rishavraj (rishavraj):

@Ratedover u shouldnt feel the same....

OpenStudy (anonymous):

Find S25 for 1/2 + 1 + 3/2 + 2 +.. how about this one, osrry I really have a hard time with sequence

rishavraj (rishavraj):

its an arithmetic series with a = 1/2 and d = 1/2...so just apply the formula S25 = (25/2)*[2a + (n - 1)d]

OpenStudy (anonymous):

OHHHH ok yeah that helps a lot thank you

rishavraj (rishavraj):

anytime!!!!!!